# Reconstruction vs. Generation: Taming Optimization Dilemma in Latent Diffusion Models

### v1

@@ 1 Introduction
However, as research has advanced, an optimization dilemma has emerged between reconstruction and generation performance in latent diffusion models [7, 13, 16]. Specifically, while increasing token feature dimension improves the tokenizer’s reconstruction accuracy, it significantly degrades the generation performance (see Fig. 1). Currently, two common strategies exist to address this issue: the first involves scaling up model parameters, as demonstrated by Stable Diffusion 3 [7], which shows that higher-dimensional tokenizers can achieve stronger generation performance with a significantly larger model capacity—however, this approach requires significantly more training compute, making it prohibitively expensive for most practical applications. The second strategy is to deliberately limit the tokenizer’s reconstruction capacity,e.g. DC-AE or W.A.L.T. [4, 13], for faster convergence of diffusion model training. Yet, this compromised reconstruction quality inherently limits the upper bound of generation performance, leading to imperfect visual details in generated results. Both approaches involve inherent trade-offs and do not fundamentally resolve the underlying optimization dilemma.
In this paper, we propose a straightforward yet effective approach to address this optimization dilemma. We draw inspiration from Auto-Regressive (AR) generation, where increasing the codebook size of discrete-valued VAEs leads to low codebook utilization [42, 47]. Through visualizing the latent space distributions across different feature dimensions (see Fig. 1), we observe that higher-dimensional tokenizers learn latent representations in a less spread-out manner, evidenced by more concentrated high-intensity areas in the distribution visualization. This analysis suggests that the optimization dilemma stems from the inherent difficulty of learning unconstrained high-dimensional latent spaces from scratch. To address this issue, we develop a vision foundation model guided optimization strategy for continuous VAEs [17] in latent diffusion models. Our key finding demonstrates that learning latent representations guided by vision foundation models significantly enhances the generation performance of high-dimensional tokenizers while preserving their original reconstruction capabilities (see Fig. 2).
@@ 2.1 Tokenizers for Visual Generation
Visual tokenizers, represented by variational autoencoders (VAEs) [17], leverage an encoder-decoder architecture to create continuous visual representations, facilitating the compression and reconstruction of visual signals. While VAEs operate in continuous space, VQVAE [39] introduces discrete representation learning through a learnable codebook for quantization. VQGAN [6] further enhances this discrete approach by incorporating adversarial training, establishing itself as a cornerstone for autoregressive (AR) generation frameworks [2]. However, these discrete approaches face a fundamental challenge: as revealed in [42], larger codebooks improve reconstruction fidelity but lead to poor codebook utilization, adversely affecting generation performance. Recent works propose different solutions to this utilization problem: MAGVIT-v2 [42] introduces Look-up Free Quantization (LFQ), while VQGAN-LC [47] leverages CLIP [32] vision features for codebook initialization, achieving near-complete codebook utilization.

### v2

@@ 1 Introduction
However, as research has advanced, an optimization dilemma has emerged between reconstruction and generation performance in latent diffusion models [7, 14, 17]. Specifically, while increasing token feature dimension improves the tokenizer’s reconstruction accuracy, it significantly degrades the generation performance (see Fig. 1). Currently, two common strategies exist to address this issue: the first involves scaling up model parameters, as demonstrated by Stable Diffusion 3 [7], which shows that higher-dimensional tokenizers can achieve stronger generation performance with a significantly larger model capacity—however, this approach requires significantly more training compute, making it prohibitively expensive for most practical applications. The second strategy is to deliberately limit the tokenizer’s reconstruction capacity, e.g. Sana [42, 4], W.A.L.T [14], for faster convergence of diffusion model training. Yet, this compromised reconstruction quality inherently limits the upper bound of generation performance, leading to imperfect visual details in generated results. Both approaches involve inherent trade-offs and do not provide an effective control to the underlying optimization dilemma.
In this paper, we propose a straightforward yet effective approach to this optimization dilemma. We draw inspiration from Auto-Regressive (AR) generation, where increasing the codebook size of discrete-valued VAEs leads to low codebook utilization [44, 49]. Through visualizing the latent space distributions across different feature dimensions (see Fig. 1), we observe that higher-dimensional tokenizers learn latent representations in a less spread-out manner, evidenced by more concentrated high-intensity areas in the distribution visualization. This analysis suggests that the optimization dilemma stems from the inherent difficulty of learning unconstrained high-dimensional latent spaces from scratch. To address this issue, we develop a vision foundation model guided optimization strategy for continuous VAEs [18] in latent diffusion models. Our key finding demonstrates that learning latent representations guided by vision foundation models significantly enhances the generation performance of high-dimensional tokenizers while preserving their original reconstruction capabilities (see Fig. 2).
@@ 2.1 Tokenizers for Visual Generation
Visual tokenizers, represented by variational autoencoders (VAEs) [18], leverage an encoder-decoder architecture to create continuous visual representations, facilitating the compression and reconstruction of visual signals. While VAEs operate in continuous space, VQVAE [40] introduces discrete representation learning through a learnable codebook for quantization. VQGAN [6] further enhances this discrete approach by incorporating adversarial training, establishing itself as a cornerstone for autoregressive (AR) generation frameworks [2]. However, these discrete approaches face a fundamental challenge: as revealed in [44], larger codebooks improve reconstruction fidelity but lead to poor codebook utilization, adversely affecting generation performance. Recent works propose different solutions to this utilization problem: SeQ-GAN [13] explores the role of semantic information in generation. MAGVIT-v2 [44] introduces Look-up Free Quantization (LFQ), while VQGAN-LC [49] leverages CLIP [33] vision features for codebook initialization, achieving near-complete codebook utilization.